### v1

# Advance Sharing Procedures for the Ramp Quantum Secret Sharing Schemes With the Highest Coding Rate

Abstract
In some quantum secret sharing schemes, it is known that some shares can be distributed to participants before a secret is given to the dealer. However, it is unclear whether some shares can be distributed before a secret is given in the ramp quantum secret sharing schemes with the highest coding rate. In this paper, we propose methods to distribute some shares before a secret is given in those schemes, then we prove that our new encoding procedures retain the correspondences between quantum secrets and quantum shares in the original schemes.
I Introduction
The study of quantum information processing gains much attention recently. One reason is that increase in the size of quantum computers [1]. The storage and communications of quantum information are still difficult experimentally, but they will probably become easier and less expensive in a future. Classical secret sharing, in which secrets and shares are both classical information, is nowadays used in practice, for example, in distributed storage systems. In a distributed storage system [2], data are stored in multiple storages, which increases the risk of data leakage. Secret sharing schemes decreases that risk with the multiple storages. It is likely that quantum secret sharing schemes play similar roles in the future quantum information era.
In this paper, we consider to share a quantum secret by quantum shares. We assume no interactions among the dealer and the participants except distribution of quantum shares from the dealer to the participants. Such a problem formulation was initiated by Cleve, Gottesman and Lo [4, 5], which is the most natural quantum counterpart of the classical secret sharing considered by Shamir [11] and Blakley [3].
The coding rate is an important parameter in secret sharing schemes. It is defined as the ratio of secret size to average of share sizes [4, 5, 10]. Higher coding rates are desirable. Another important property of secret sharing schemes is the access structure, which consists of three families of qualified sets, forbidden sets, and intermediate sets [10]. A set $S$ of shares is called qualified (resp.  forbidden) if $S$ allows reconstruction of the secret (resp. has no information about the secret). A set $S$ of shares is called intermediate if it is neither qualified nor forbidden. If a secret sharing scheme has no intermediate set, it is called perfect. The coding rate cannot be greater than $1$ if it is perfect [5]. Ogawa et al. proposed ramp quantum secret sharing schemes, which enable coding rates greater than $1$, at the cost of allowing intermediate sets. In this paper, we consider $q$-dimensional quantum system $\mathcal{H}_{q}$ and call it a qudit, where $q$ is a prime power. A $(k,L,n)$ ramp quantum secret sharing scheme encodes $L$-qudit quantum secret into $n$ shares each of which is 1 qudit in $\mathcal{H}_{q}$. In a $(k,L,n)$ scheme, a set $S$ is qualified if $|S|\geq k$, forbidden if $|S|\leq k-L$, and intermediate otherwise (i.e. $k-L+1\leq|S|\leq k-1$). Among $(k,L,n)$ schemes, Ogawa et al.’s one has the highest possible coding rate $L$, and firstly we consider Ogawa et al.’s ramp quantum secret sharing in Section II. The scheme by Cleve, Gottesman and Lo [4] is a special case of Ogawa et al.’s corresponding to the case $L=1$.
As mentioned before, an intermediate set has nonzero information about the secret in a ramp scheme. In classical secret sharing scheme, when the secret $\vec{s}=(s_{1}$, …, $s_{L})$, Iwamoto and Yamamoto [6] showed an explicit example of ramp schemes in which a component $s_{i}$ in the secret can be reconstructed from an intermediate set $S$. In order to prevent such information leakage, Yamamoto [13] defined the strong security for $(k,L,n)$ classical ramp secret sharing scheme, in which any set $S$ of shares has no information about part $(s_{i}:i\in T)$ of the secret if $|S|+|T|\leq k$. In the context of ramp quantum secret sharing, Zhang and Matsumoto [14] showed an explicit example from Ogawa et al.’s scheme in which an intermediate set leaks part of the quantum secret similarly to the classical case in this paragraph, introduced a strong security definition into ramp quantum secret sharing, and explicitly construct a $(k,L,n)$ strongly secure ramp quantum secret sharing scheme whose coding rate is as high as Ogawa et al.’s scheme [10]. On the other hand, the Zhang-Matsumoto scheme has a more stringent condition on the number $n$ of participants, that is, $n\leq q-L$, while Ogawa et al.’s scheme has a milder condition $n\leq q-1$. The condition $n\leq q-1$ also exists in Shamir’s scheme [11, 3], but in the classical case $q$ can be made almost arbitrarily large. On the other hand, because the quantum system $\mathcal{H}_{q}$ cannot be chosen freely, the symbol size $q$ cannot be adjusted in the quantum case as easy as the classical case. Therefore, both Ogawa et al.’s scheme and Zhang-Matsumoto scheme are practically useful depending on applications.
It is sometimes convenient to distribute shares before a secret is given to the dealer. One example of such situations was discussed in [9], which considered sharing a classical secret by quantum shares. Such distribution of shares before a given secret is named “advance sharing” [9]. Advance sharing is a trivial problem when both secret and shares are classical. For example, encoding 1-bit secret $s$ into 2 shares $(x,s+x)$ provides a $(1,2)$ perfect secret sharing scheme and the first share $x$ is clearly advance shareable, where $x$ is a random bit. On the other hand, when shares are quantum, it is nontrivial to realize advance sharing. In this paper, we focus on the case of quantum secrets. The first quantum advance sharing was realized in Lie and Jeong [7], which enabled advance sharing for some specific schemes. On the other hand, any quantum error-correcting code and its erasure decoding algorithm can be used as a quantum secret sharing scheme [4, 5], and Shibata and Matsumoto [12] clarified how to realize advance sharing with quantum secret sharing scheme constructed from a $p$-adic quantum stabilizer code, where $p$ is a prime number. Based on [12], Matsumoto and Masumori [8] showed how to realize advance sharing with a limited special case of [10], where the dimension $q$ of $\mathcal{H}_{q}$ had to be a prime number. Because the general method [12] for advance sharing cannot handle stabilizer codes over $\mathcal{H}_{q}$ with non-prime $q$, the previous proposal [8] of advance sharing for [10] cannot be immediately extended to the general case of [10].
When we consider advance sharing for the strongly secure scheme [14], another limitation of [12] appears. The strong security property in [14] is realized by a careful correspondence between quantum secrets and quantum shares. On the other hand, the general method [12] for advance sharing destroys the correspondence between quantum secrets and quantum shares when it is applied to a ramp quantum secret sharing scheme. The general method [12] seems almost impossible to be applied for realizing advance sharing with the strongly secure scheme [14].
In this paper, we propose new encoding procedures of quantum secrets into quantum shares for realizing advance sharing with Ogawa et al.’s scheme [10] and Zhang-Matsumoto scheme [14]. Both procedures retain the correspondence between quantum secrets and quantum shares in the original schemes. Therefore, all properties in the original schemes, such as coding rates, access structures, strong security, etc., remain the same as the originals. The proposed new encoding procedures add extra useful functionalities to the original ramp quantum schemes [10, 14].
This paper is organized as follows: In Section II, we review necessary contents from [10], a new encoding procedure is given for [10], and we prove that the correspondence between secrets and shares is retained from [10]. Section III has a similar structure to Section II. In Section III, we review [14], propose a new encoding procedure for [14], then we prove that the same correspondence between secrets and shares. Throughout Sections II and III, illustrating examples of the original and the proposed encoding are included. In Section IV, we give concluding remarks and future research agendas.
II Advance sharing with quantum ramp secret sharing scheme in [10]
II-A Ogawa et al.’s scheme
Let $q$ be a prime power, $\mathbf{F}_{q}$ be the finite field with $q$ elements, and $\{|i\rangle:i\in\mathbf{F}_{q}\}$ be an orthonormal basis of $\mathcal{H}_{q}$. For $\vec{s}=(s_{1}$, …, $s_{L})\in\mathbf{F}_{q}^{L}$ we define
$|\vec{s}\rangle=|s_{1}\rangle\otimes\cdots\otimes|s_{L}\rangle\in\mathcal{H}_{q}^{\otimes L}.$
Let $n$ be the number of shares/participants, and we assume $n<q$. Let $\alpha_{1}$, …, $\alpha_{n}$ be distinct nonzero [^1] Ogawa et al. [10] allowed zero but it was a mistake. elements in $\mathbf{F}_{q}$. We will review the construction of a $(k,L,n)$ ramp quantum scheme. For $\vec{c}=(c_{1}$, …, $c_{k})\in\mathbf{F}_{q}^{k}$, we define the polynomial $f_{\vec{c}}(x)$ by
$f_{\vec{c}}(x)=c_{1}+c_{2}x+\cdots+c_{k}x^{k-1}.$ (1)
A quantum secret $|\vec{s}\rangle\in\mathcal{H}_{q}^{\otimes L}$ is encoded to
$\frac{1}{\sqrt{q^{k-L}}}\sum_{\vec{c}\in D(\vec{s})}|f_{\vec{c}}(\alpha_{1})\rangle\otimes|f_{\vec{c}}(\alpha_{2})\rangle\otimes\cdots\otimes|f_{\vec{c}}(\alpha_{n})\rangle,$ (2)
where $D(\vec{s})$ is the set of vectors $\vec{c}\in\mathbf{F}_{q}^{k}$ whose leftmost $L$ components are the same as those in $\vec{s}$. The quantum state in Eq.  (2) consists of $n$ qudits. The $i$-th qudit in Eq.  (2) is distributed to the $i$-th participant.
II-B Example of the original encoding
Let $p=2$, $q=4$, $n=3$, $k=2$, $L=1$, and $\alpha_{i}=\alpha^{i}\in\mathbf{F}_{4}$, where $\alpha$ is a primitive element of $\mathbf{F}_{4}$. The standard encoding procedure in Eq. (2) encodes a quantum secret $|\vec{s}\rangle$ for $\vec{s}=(s)\in\mathbf{F}_{4}^{1}$ into $\frac{1}{\sqrt{4}}\sum_{\vec{c}\in D(\vec{s})}|f_{\vec{c}}(\alpha_{1})\rangle\otimes|f_{\vec{c}}(\alpha_{2})\rangle\otimes|f_{\vec{c}}(\alpha_{3})\rangle=\frac{1}{\sqrt{4}}(|s$, $s$, $s\rangle+|s+\alpha$, $s+\alpha^{2}$, $s+1\rangle+|s+\alpha^{2}$, $s+1$, $s+\alpha\rangle+|s+1$, $s+\alpha$, $s+\alpha^{2}\rangle).$ As every qudit depends on the quantum secret $|\vec{s}\rangle$, from the encoding procedure in [10], it seems unclear (to the authors) how one share is advance shareable.
II-C Proposed advance sharing procedure for [10]
We will propose an advance sharing procedure which retains the correspondence between a quantum secret and quantum shares given in Eq. (2), where the $(k,L,n)$ quantum ramp scheme was considered.
We will introduce an elementary lemma in linear algebra and polynomials.
Lemma 1
Let $\mathcal{P}_{m}$ be the set of univariate polynomials $f(x)$ over $\mathbf{F}_{q}$ with $\deg(f)<m$. Consider the evaluation map $\mathrm{ev}(f)=(f(\alpha_{1})$, …, $f(\alpha_{n}))$, where $\alpha_{1}$, …, $\alpha_{n}$ are pairwise distinct as before while $\alpha_{1}$, …, $\alpha_{n}$ may contain $0\in\mathbf{F}_{q}$. The map $\mathrm{ev}$ is a linear map from $\mathcal{P}_{m}$. If $m<n$ then $\mathrm{ev}$ is injective, and if $m\geq n-1$ then $\mathrm{ev}$ is surjective.
Proof:
Suppose that $m<n$. If $\mathrm{ev}(f)=\vec{0}$ then $f(\alpha_{1})=\cdots=f(\alpha_{n})=0$. Since $\deg(f)\leq m<n$, this means that $f(x)$ is the zero polynomial. We have seen that $\ker(\mathrm{ev})=\{0\}$ when $m<n$, which means $\mathrm{ev}$ is an injective linear map.
By the previous discussion, $\mathrm{ev}$ is injective on $\mathcal{P}_{n-1}$. Since $\dim\mathcal{P}_{n-1}=n$, the image $\mathrm{ev}(\mathcal{P}_{n-1})$ of $\mathcal{P}_{n-1}$ under the map $\mathrm{ev}$ is $\mathbf{F}_{q}^{n}$. This means that $\mathrm{ev}$ is surjective on $\mathcal{P}_{m}$ if $m\geq n-1$. ∎
We consider to distribute $k-L$ shares before $|\vec{s}\rangle$ is given to the dealer. By reassigning indices, we may assume that the dealer wants to advance share the 1st to the $(k-L)$-th shares. Let
$|\Psi\rangle=\frac{1}{q^{k-L}}\sum_{\vec{r}\in\mathbf{F}_{q}^{k-L}}|\vec{r}\rangle\otimes|\vec{r}\rangle,$ (3)
which consists of $2(k-L)$ qudits.
Suppose that the quantum secret is $|\vec{s}\rangle=|s_{1}\rangle\otimes\cdots\otimes|s_{L}\rangle\in\mathcal{H}_{q}^{\otimes L}$ with $\vec{s}=(s_{1}$, …, $s_{L})\in\mathbf{F}_{q}^{L}$. For $\vec{r}\in\mathbf{F}_{q}^{k-L}$ and $\vec{s}$, let $g_{\vec{r},\vec{s}}(x)$ be a univariate polynomial such that $g_{\vec{r},\vec{s}}(x)\in D(\vec{s})$ and $g_{\vec{r},\vec{s}}(x)(\alpha_{i})=r_{i}$ for $i=1$, …, $k-L$.
Proposition 2
For given $\vec{r}$ and $\vec{s}$, the polynomial $g_{\vec{r},\vec{s}}(x)$ exists uniquely.
Proof:
Let $f_{\vec{c}}(x)\in D(\vec{s})$ as defined in Eqs. (1) and (2). Recall that the coefficients $c_{1}$, …, $c_{L}$ are known as $c_{1}=s_{1}$, …, $c_{L}=s_{L}$ by the condition $f_{\vec{c}}(x)\in D(\vec{s})$. We will regard the coefficients $c_{L+1}$, …, $c_{k}$ in $f_{\vec{c}}(x)$ as $k-L$ unknowns in a system of linear equations. For $i=1$, …, $k-L$, the condition $g_{\vec{r},\vec{s}}(x)(\alpha_{i})=r_{i}$ can be written as
$c_{L+1}+c_{L+1}\alpha_{i}+\cdots+c_{k}\alpha_{i}^{k-L-1}=\frac{r_{i}-s_{1}-s_{2}\alpha_{i}-\cdots-s_{L}\alpha_{i}^{L-1}}{\alpha_{i}^{L}}.$
The polynomial $f(x)=c_{L+1}+\cdots+c_{k}x^{k-L-1}\in\mathcal{P}_{k-L}$ corresponds to a solution of the system of linear equations if $\mathrm{ev}(f)=((r_{1}-s_{1}-s_{2}\alpha_{1}-\cdots-s_{L}\alpha_{1}^{L-1})/\alpha_{1}^{L}$, $(r_{2}-s_{1}-s_{2}\alpha_{2}-\cdots-s_{L}\alpha_{2}^{L-1})/\alpha_{2}^{L}$, …, $(r_{k-L}-s_{1}-s_{2}\alpha_{k-L}-\cdots-s_{L}\alpha_{k-L}^{L-1})/\alpha_{k-L}^{L})$, where the map $\mathrm{ev}$ is from $\mathcal{P}_{k-L}$ to $\mathbf{F}_{q}^{k-L}$. By Lemma 1 we can see that the solution of the system of linear equations considered here exists uniquely, which ensures unique existence of $g_{\vec{r},\vec{s}}(x)$ for given $\vec{r}$ and $\vec{s}$. ∎
By Proposition 2, we can define a unitary map $U_{\mathrm{enc}}$ from $\mathcal{H}_{q}^{\otimes k}$ to a subspace of $\mathcal{H}_{q}^{\otimes(n-k+L)}$ sending a quantum state $|\vec{r}\rangle|\vec{s}\rangle$ to $|g_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|g_{\vec{r},\vec{s}}(\alpha_{n})\rangle$.
The proposed procedure for advance sharing is as follows:
  1. 1.
    The dealer distributes $k-L$ qudits in the left half of Eq. (3).
  2. 2.
    After the quantum secret $|\vec{s}\rangle$ is given, the dealer applies $U_{\mathrm{enc}}$ to the remaining half of Eq. (3) and $|\vec{s}\rangle$. Then the quantum state of all the $n$ shares becomes
    $\frac{1}{q^{k-L}}\sum_{\vec{r}\in\mathbf{F}_{q}^{k-L}}|\vec{r}\rangle|g_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|g_{\vec{r},\vec{s}}(\alpha_{n})\rangle\in\mathcal{H}_{q}^{\otimes n}.$ (4)
Remark 3
When $k-L-1$ or a fewer shares are advance shared, the dealer can simply keep some shares in the advance sharing phase in our proposal. On the other hand, it is impossible to advance share $k-L+1$ or more shares. When a set $J$ of shares is not forbidden, the shares in $J$ depend on quantum secrets [10, Theorem 2]. Thus, in order for a set $J$ of shares to be advance shareable, $J$ must be a forbidden set. In Ogawa et al.’s $(k,L,n)$ scheme, any $k-L+1$ or more shares cannot form a forbidden set, and cannot be advance shareable. Our proposal makes advance shareable sets as large as possible.
II-D Proof of correctness
In this subsection, we will prove that the proposed procedure gives the same quantum state of $n$ shares as the original [10] for a given quantum secret $|\vec{s}\rangle$. In Eq. (2), the set of indices appearing in the quantum state is
$\{(f_{\vec{c}}(\alpha_{1}),\ldots,f_{\vec{c}}(\alpha_{n})):\vec{c}\in D(\vec{s})\}.$ (5)
In Eq. (4), the set of indices appearing in the quantum state is
$\{(r_{1},\ldots,r_{k-L},g_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,g_{\vec{r},\vec{s}}(\alpha_{n})):\vec{r}\in\mathbf{F}_{q}^{k-L}\}.$ (6)
Theorem 4
For a fixed quantum secret $|\vec{s}\rangle$ for $\vec{s}\in\mathbf{F}_{q}^{L}$, the proposed procedure gives the same quantum state of shares as the original [10].
Proof:
In order to prove the theorem, it is enough to show that the sets (5) and (6) are the same. For fixed $\vec{s}\in\mathbf{F}_{q}^{L}$ and $\vec{r}\in\mathbf{F}_{q}^{k-L}$, by Proposition 2 there exists a unique polynomial $f_{\vec{c}}$ with $\vec{c}\in D(\vec{s})$. This correspondence gives a bijectition sending $\vec{r}\in\mathbf{F}_{q}^{k-L}$ to $f_{\vec{c}}$ with $\vec{c}\in D(\vec{s})$. By the definition of $g_{\vec{r},\vec{s}}$, we have
$(r_{1},\ldots,r_{k-L},g_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,g_{\vec{r},\vec{s}}(\alpha_{n}))$
$=$ $(g_{\vec{r},\vec{s}}(\alpha_{1}),\ldots,g_{\vec{r},\vec{s}}(\alpha_{n}))$
$=$ $(f_{\vec{c}}(\alpha_{1}),\ldots,f_{\vec{c}}(\alpha_{n})),$
which shows the theorem, where $\vec{c}$ is chosen according to $\vec{r}$. ∎
II-E Example of the proposed encoding
We reuse definitions from Section II-B. Let $p=2$, $q=4$, $n=3$, $k=2$, $L=1$, and $\alpha_{i}=\alpha^{i}\in\mathbf{F}_{4}$, where $\alpha$ is a primitive element of $\mathbf{F}_{4}$. Note that since $p\neq q$, this case cannot be handled in the previous research [8]. In our proposal, the dealer prepares 2-qudit entangled state $\frac{1}{2}\sum_{r\in\mathbf{F}_{4}}|r\rangle|r\rangle$, and send 1 qudit in it. Then the dealer is given a quantum secret $|\vec{s}\rangle$ for $\vec{s}=(s)\in\mathbf{F}_{4}^{1}$. Then we have
$g_{r,s}(x)=\frac{r-s}{\alpha_{1}}x+s.$
The unitary map $U_{\mathrm{enc}}$ sends $|r\rangle|s\rangle$ to $|\frac{r-s}{\alpha_{1}}\alpha_{2}+s\rangle|\frac{r-s}{\alpha_{1}}\alpha_{3}+s\rangle$. Now it is clear that the left most share can be distributed before the secret $|s\rangle$ is given, in contrast to Section II-B.
III Advance sharing with quantum ramp secret sharing scheme in [14]
III-A Zhang and Matsumoto’s scheme
As mentioned before, Zhang and Matsumoto [14] also proposed a $(k,L,n)$ ramp quantum secret sharing scheme. Their proposal has strong security in contrast to [10], and the maximum possible number of participants was smaller, that is, $n\leq q-L$. As in Section II, $\alpha_{1}$, …, $\alpha_{n}$ are pairwise distinct elements in $\mathbf{F}_{q}$. We choose extra elements $\beta_{1}$, …, $\beta_{L}$ such that all $\alpha_{1}$, …, $\alpha_{n}$ and $\beta_{1}$, …, $\beta_{L}$ are pairwise distinct. In [14] any of $\alpha_{1}$, …, $\alpha_{n}$ and $\beta_{1}$, …, $\beta_{L}$ can be $0\in\mathbf{F}_{q}$. For $\vec{s}\in\mathbf{F}_{q}^{L}$, we define
$D_{\mathrm{ZM}}(\vec{s})=\{\vec{c}\in\mathbf{F}_{q}^{k}:f_{\vec{c}}(\beta_{i})=s_{i}\mbox{ for }i=1,\ldots,L\},$ (7)
where $f_{\vec{c}}(x)$ is as defined in Eq. (1). For a given quantum secret $|\vec{s}\rangle$, in [14] the quantum state of $n$ shares are obtained as
$\frac{1}{\sqrt{q^{k-L}}}\sum_{\vec{c}\in D_{\mathrm{ZM}}(\vec{s})}|f_{\vec{c}}(\alpha_{1})\rangle\otimes|f_{\vec{c}}(\alpha_{2})\rangle\otimes\cdots\otimes|f_{\vec{c}}(\alpha_{n})\rangle,$ (8)
III-B Example of the original encoding
Let $p=q=7$, $n=4$, $k=3$, $L=2$, $(\alpha_{1}$, …, $\alpha_{4})=(6$, $2$, $4$, 5)5)5 ) and $(\beta_{1}$, $\beta_{2})=(1$, 3)3)3 ). This example is the same as [14, Example 2]. For $\vec{s}=(s_{1}$, $s_{2})$, $D_{\mathrm{ZM}}(\vec{s})$ contains $(c_{1}$, …, $c_{3})$ if and only if
$\left\{\begin{array}[]{rcl}c_{1}+c_{2}+c_{3}&=&s_{1},\\ c_{1}+3c_{2}+2c_{3}&=&s_{2}.\end{array}\right.$
The solution of the above system of linear equations for $(c_{1}$, …, $c_{3})$ is $D_{\mathrm{ZM}}(\vec{s})=\{(5s_{1}+3s_{2}+3c_{3},3s_{1}+4s_{2}+3c_{3},c_{3}):c_{3}\in\mathbf{F}_{7}\}$, which is the same [14, Eq. (3)]. For a given quantum state $|\vec{s}\rangle=|s_{1}\rangle\otimes|s_{2}\rangle$, the quantum state of $4$ shares is (the normalizing constant $1/\sqrt{7}$ is omitted) $\sum_{c_{3}=0}^{6}\otimes_{i=1}^{4}|\alpha_{i}^{2}c_{3}+\alpha_{i}(3s_{1}+4s_{2}+3c_{3})+(5s_{1}+3s_{2}+3c_{3})\rangle=|2s_{1}+6s_{2}\rangle\otimes|4s_{1}+4s_{2}\rangle\otimes|3s_{1}+5s_{2}\rangle\otimes|6s_{1}+2s_{2}\rangle+\cdots$, which is equivalent to [14, Eq. (5)]. Since the quantum state of every qudit depends on $s_{1}$ and $s_{2}$, from the encoding procedure in [14], it seems unclear (to the authors) how one share is advance shareable.
III-C Proposed advance sharing procedure for [14]
We consider to distribute $k-L$ shares before $|\vec{s}\rangle$ is given to the dealer. By reassigning indices, we may assume that the dealer wants to advance share the 1st to the $(k-L)$-th shares. Let $|\Psi\rangle$ as defined in Eq. (3).
Suppose that the quantum secret is $|\vec{s}\rangle=|s_{1}\rangle\otimes\cdots\otimes|s_{L}\rangle\in\mathcal{H}_{q}^{\otimes L}$ with $\vec{s}=(s_{1}$, …, $s_{L})\in\mathbf{F}_{q}^{L}$. For $\vec{r}\in\mathbf{F}_{q}^{k-L}$ and $\vec{s}$, let $h_{\vec{r},\vec{s}}(x)$ be a univariate polynomial such that $h_{\vec{r},\vec{s}}(x)\in D_{\mathrm{ZM}}(\vec{s})$ and $h_{\vec{r},\vec{s}}(x)(\alpha_{i})=r_{i}$ for $i=1$, …, $k-L$.
Proposition 5
For given $\vec{r}$ and $\vec{s}$, the polynomial $h_{\vec{r},\vec{s}}(x)$ exists uniquely.
Proof:
Let $f(x)\in\mathcal{P}_{k}$ as defined in Lemma 1. The condition $f(x)\in D_{\mathrm{ZM}}(\vec{s})$ means that $f(\beta_{i})=s_{i}$ for $i=1$, …, $L$. Together with the conditions $f(x)(\alpha_{i})=r_{i}$ for $i=1$, …, $k-L$, the required $h_{\vec{r},\vec{s}}(x)$ is a $f(x)$ such that $\mathrm{ev}^{\prime}(f)=(s_{1}$, …, $s_{L}$, $r_{1}$, …, $r_{k-L})$, where $\mathrm{ev}^{\prime}(f)=(f(\beta_{1})$, …, $f(\beta_{L})$, $f(\alpha_{1})$, …, $f(\alpha_{k-L}))$. By Lemma 1, we can see that such a polynomial $f(x)$ exists uniquely, which is $h_{\vec{r},\vec{s}}(x)$. ∎
By Proposition 5, we can define a unitary map $U_{\mathrm{ZM,enc}}$ from $\mathcal{H}_{q}^{\otimes k}$ to $\mathcal{H}_{q}^{\otimes(n-k+L)}$ sending a quantum state $|\vec{r}\rangle|\vec{s}\rangle$ to $|h_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|h_{\vec{r},\vec{s}}(\alpha_{n})\rangle$.
The proposed procedure for advance sharing is as follows:
  1. 1.
    The dealer distribute $k-L$ qudits in the left half of Eq. (3).
  2. 2.
    After the quantum secret $|\vec{s}\rangle$ is given, the dealer applies $U_{\mathrm{ZM,enc}}$ to the remaining half of Eq. (3) and $|\vec{s}\rangle$. Then the quantum state of all the $n$ shares becomes
    $\frac{1}{q^{k-L}}\sum_{\vec{r}\in\mathbf{F}_{q}^{k-L}}|\vec{r}\rangle|h_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|h_{\vec{r},\vec{s}}(\alpha_{n})\rangle\in\mathcal{H}_{q}^{\otimes n}.$ (9)
Remark 6
When $k-L-1$ or a fewer shares are advance shared, the dealer can simply keep some shares in the advance sharing phase in our proposal as in Remark 3. On the other hand, as in Remark 3, it is impossible to advance share $k-L+1$ or more shares, because in Zhang and Matsumoto’s $(k,L,n)$ scheme [14], any $k-L+1$ or more shares cannot form a forbidden set. Therefore, similarly to Remark 3, our proposal makes advance shareable sets as large as possible.
III-D Proof of correctness
In this subsection, we will prove that the proposed procedure gives the same quantum state of shares as the original [14] for a given quantum secret $|\vec{s}\rangle$. In Eq. (8), the set of indices appearing in the quantum state is
$\{(f_{\vec{c}}(\alpha_{1}),\ldots,f_{\vec{c}}(\alpha_{n})):\vec{c}\in D_{\mathrm{ZM}}(\vec{s})\}.$ (10)
In Eq. (9), the set of indices appearing in the quantum state is
$\{(r_{1},\ldots,r_{k-L},h_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,h_{\vec{r},\vec{s}}(\alpha_{n}):\vec{r}\in\mathbf{F}_{q}^{k-L}\}.$ (11)
Theorem 7
For a fixed quantum secret $|\vec{s}\rangle$ for $\vec{s}\in\mathbf{F}_{q}^{L}$, the proposed procedure gives the same quantum state of shares as the original [14].
Proof:
In order to prove the theorem, it is enough to show that the sets (10) and (11) are the same. For fixed $\vec{s}\in\mathbf{F}_{q}^{L}$ and $\vec{r}\in\mathbf{F}_{q}^{k-L}$, by Proposition 5 there exists a unique polynomial $f_{\vec{c}}(x)$ with $\vec{c}\in D_{\mathrm{ZM}}(\vec{s})$. This correspondence gives a bijectition between $\vec{r}\in\mathbf{F}_{q}^{k-L}$ to $f_{\vec{c}}(x)$ with $\vec{c}\in D(\vec{s})$. By the definition of $h_{\vec{r},\vec{s}}$, we have
$(r_{1},\ldots,r_{k-L},h_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,h_{\vec{r},\vec{s}}(\alpha_{n}))$
$=$ $(h_{\vec{r},\vec{s}}(\alpha_{1}),\ldots,h_{\vec{r},\vec{s}}(\alpha_{n}))$
$=$ $(f_{\vec{c}}(\alpha_{1}),\ldots,f_{\vec{c}}(\alpha_{n})),$
which shows the theorem, where $\vec{c}$ is chosen according to $\vec{r}$. ∎
III-E Example of the proposed encoding
We reuse definitions from Section III-B. Let $p=q=7$, $n=4$, $k=3$, $L=2$, $(\alpha_{1}$, …, $\alpha_{4})=(6$, $2$, $4$, 5)5)5 ) and $(\beta_{1}$, $\beta_{2})=(1$, 3)3)3 ). In our proposal, the dealer prepares 2-qudit entangled state $\frac{1}{\sqrt{7}}\sum_{r\in\mathbf{F}_{7}}|r\rangle|r\rangle$, and send 1 qudit in it. Then the dealer is given a quantum secret $|\vec{s}\rangle$ for $\vec{s}=(s_{1},s_{2})\in\mathbf{F}_{7}^{2}$. Then we have
$h_{r,s_{1},s_{2}}(x)=(r-2s_{1}+s_{2})x^{2}+(4s_{1}-4r)x+(3r-s_{1}-s_{2}),$
because $h_{r,s_{1},s_{2}}(x)$ satisfies $h_{r,s_{1},s_{2}}(\alpha_{1})=h_{r,s_{1},s_{2}}(6)=r$ , $h_{r,s_{1},s_{2}}(\beta_{1})=h_{r,s_{1},s_{2}}(1)=s_{1}$, and $h_{r,s_{1},s_{2}}(\beta_{2})=h_{r,s_{1},s_{2}}(3)=s_{2}$. The unitary map $U_{\mathrm{ZM,enc}}$ sends $|r\rangle|s_{1}\rangle|s_{2}\rangle$ to $|h_{r,s_{1},s_{2}}(\alpha_{2})\rangle|h_{r,s_{1},s_{2}}(\alpha_{3})\rangle|h_{r,s_{1},s_{2}}(\alpha_{3})\rangle$. Now it is clear that the left most share can be distributed before the secret $|\vec{s}\rangle$ is given, in contrast to Section III-B.
IV Concluding remarks
In this paper, we showed how to make shares in Ogawa et al.’s [10] and Zhang-Matsumoto [14] ramp quantum secret sharing scheme advance shareable, and also proved that the new proposed encoding procedures retain all properties, such as coding rates, access structures, strong security from the original schemes. We also proved that the advance shareable sets are made as large as possible in the proposed encoding procedures.
Similarly to [12], there could be a general method to construct advance sharing procedure for any stabilizer-based quantum secret sharing, while retaining the correspondence between secrets and shares. Its investigation could be a future research agenda.
Acknowledgment
This research is in part supported by the JSPS grant No. 23K10980.