### v2

Abstract
In some quantum secret sharing schemes, it is known that some shares can be distributed to participants before a secret is given to the dealer. However, it is unclear whether some shares can be distributed before a secret is given in the ramp quantum secret sharing schemes with the highest coding rate. In this paper, we propose procedures to distribute some shares before a secret is given in those schemes. The new procedures enhances applicability of the secret sharing schemes to wider scenarios as some participants can be unavailable when the dealer obtains the quantum secret. Then we prove that our new encoding procedures retain the correspondences between quantum secrets and quantum shares in the original schemes, which ensures the highest coding rates of the original schemes are also retained.
Index Terms:
advance sharing, quantum secret sharing, ramp secret sharing
[…]
=-15pt
I Introduction
The study of quantum information processing gains much attention recently. One reason is the increase in the size of quantum computers [1]. The storage and communications of quantum information are still difficult experimentally, but they will probably become easier and less expensive in a future. Classical secret sharing, in which secrets and shares are both classical information, is nowadays used in practice, for example, in distributed storage systems. In a distributed storage system [2], data are stored in multiple storages, which increases the risk of data leakage. Secret sharing schemes decreases that risk with the multiple storages. It is likely that quantum secret sharing schemes play similar roles in the future quantum information era.
TABLE I: Comparisons of the original and the proposed encoding procedures
Ogawa et al. [7] Zhang-Matsumoto [10] Section II Section III
dimension of a quantum share $q$ $q$ $q$ $q$
number of symbols in quantum secrets $L$ $L$ $L$ $L$
maximum number of participants $q-1$ $q-L$ $q-1$ $q-L$
coding rate $L$ $L$ $L$ $L$
minimum number of participants to reconstruct secrets $k$ $k$ $k$ $k$
maximum number of participants with no information about secrets $k-L$ $k-L$ $k-L$ $k-L$
strong security No Yes No Yes
maximum number of shares distributed in advance 00 00 $k-L$ $k-L$
In this paper, we consider to share a quantum secret by quantum shares. We assume no interactions among the dealer and the participants except distribution of quantum shares from the dealer to the participants. Such a problem formulation was initiated by Cleve, Gottesman and Lo [3, 4], which is the most natural quantum counterpart of the classical secret sharing considered by Shamir [5] and Blakley [6].
The coding rate is an important parameter in secret sharing schemes. It is defined as the ratio of secret size to average of share sizes [3, 4, 7]. Higher coding rates are desirable. Another important property of secret sharing schemes is the access structure, which consists of three families of qualified sets, forbidden sets, and intermediate sets [7]. A set $S$ of shares is called qualified (resp.  forbidden) if $S$ allows reconstruction of the secret (resp. has no information about the secret). A set $S$ of shares is called intermediate if it is neither qualified nor forbidden. If a secret sharing scheme has no intermediate set, it is called perfect. The coding rate cannot be greater than $1$ if it is perfect [4]. Ogawa et al. proposed ramp quantum secret sharing schemes, which enable coding rates greater than $1$, at the cost of allowing intermediate sets. In this paper, we consider $q$-dimensional quantum system $\mathcal{H}_{q}$ and call it a qudit, where $q$ is a prime power. A $(k,L,n)$ ramp quantum secret sharing scheme encodes $L$-qudit quantum secret into $n$ shares each of which is 1 qudit in $\mathcal{H}_{q}$. The parameter $k$ decides the access structure mentioned earlier, as follows. In a $(k,L,n)$ scheme, a set $S$ is qualified if $|S|\geq k$, forbidden if $|S|\leq k-L$, and intermediate otherwise (i.e. $k-L+1\leq|S|\leq k-1$). Among $(k,L,n)$ schemes, Ogawa et al.’s one has the highest possible coding rate $L$, and firstly we consider Ogawa et al.’s ramp quantum secret sharing in Section II. The scheme by Cleve, Gottesman and Lo [3] is a special case of Ogawa et al.’s corresponding to the case $L=1$.
As mentioned before, an intermediate set has nonzero information about the secret in a ramp scheme. In classical secret sharing scheme, when the secret $\vec{s}=(s_{1}$, …, $s_{L})$, Iwamoto and Yamamoto [8] showed an explicit example of ramp schemes in which a component $s_{i}$ in the secret can be reconstructed from an intermediate set $S$. In order to prevent such information leakage, Yamamoto [9] defined the strong security for $(k,L,n)$ classical ramp secret sharing scheme, in which any set $S$ of shares has no information about part $(s_{i}:i\in T)$ of the secret if $|S|+|T|\leq k$. In the context of ramp quantum secret sharing, Zhang and Matsumoto [10] showed an explicit example from Ogawa et al.’s scheme in which an intermediate set leaks part of the quantum secret similarly to the classical case in this paragraph. They also introduced a strong security definition into ramp quantum secret sharing, and explicitly construct a $(k,L,n)$ strongly secure ramp quantum secret sharing scheme whose coding rate $L$ is as high as Ogawa et al.’s scheme [7]. On the other hand, the Zhang-Matsumoto scheme has a more stringent condition on the number $n$ of participants, that is, $n\leq q-L$, while Ogawa et al.’s scheme has a milder condition $n\leq q-1$. The condition $n\leq q-1$ also exists in Shamir’s scheme [5, 6], but in the classical case $q$ can be made almost arbitrarily large. On the other hand, because the quantum system $\mathcal{H}_{q}$ cannot be chosen freely, the symbol size $q$ cannot be adjusted in the quantum case as easily as the classical case. Therefore, both Ogawa et al.’s scheme and Zhang-Matsumoto scheme are practically useful depending on applications.
It is sometimes convenient to distribute shares before a secret is given to the dealer. One example of such situations was discussed in [11], which considered sharing a classical secret by quantum shares. Such distribution of shares before a given secret is named “advance sharing” [11]. Advance sharing is a trivial problem when both secret and shares are classical. For example, encoding 1-bit secret $s$ into 2 shares $(x,s+x)$ provides a $(1,2)$ perfect secret sharing scheme and the first share $x$ is clearly advance shareable, where $x$ is a random bit. Advance sharing enhances applicability of secret sharing schemes to wider scenarios.
On the other hand, when secrets are quantum, it is nontrivial to realize advance sharing. In this paper, we focus on the case of quantum secrets. The first quantum advance sharing was realized in Lie and Jeong [12], which enabled advance sharing for some specific schemes. On the other hand, any quantum error-correcting code and its erasure decoding algorithm can be used as a quantum secret sharing scheme [3, 4], and Shibata and Matsumoto [13] clarified how to realize advance sharing with quantum secret sharing scheme constructed from a $p$-adic quantum stabilizer code, where $p$ is a prime number. Based on [13], Masumori and Matsumoto [14] showed how to realize advance sharing with a limited special case of Ogawa et al. [7], where the dimension $q$ of $\mathcal{H}_{q}$ was restricted to a prime number. Because the general method [13] for advance sharing cannot handle stabilizer codes over $\mathcal{H}_{q}$ with non-prime $q$, the previous proposal [14] of advance sharing for Ogawa et al. [7] cannot be immediately extended to the general case of Ogawa et al. [7].
When we consider advance sharing for the strongly secure scheme [10], another limitation of [13] appears. The strong security property in [10] is realized by a careful correspondence between quantum secrets and quantum shares. On the other hand, the general method [13] for advance sharing destroys the correspondence between quantum secrets and quantum shares when it is applied to a ramp quantum secret sharing scheme. The general method [13] seems almost impossible to be applied for realizing advance sharing with the strongly secure scheme [10].
In this paper, we propose new encoding procedures of quantum secrets into quantum shares for realizing advance sharing with Ogawa et al.’s scheme [7] and Zhang-Matsumoto scheme [10]. The new procedures enhance applicability of the scheme proposed in [7, 10] to wider scenarios in which some participants are unavailable when the dealer obtains quantum secrets to be shared. Both procedures retain the correspondence between quantum secrets and quantum shares in the original schemes. Therefore, all properties in the original schemes, such as coding rates, access structures, strong security, etc., remain the same as the originals. In particular, the highest coding rates are retained from the original schemes [7, 10]. The proposed new encoding procedures add extra useful functionalities to the original ramp quantum schemes [7, 10]. The differences among the original schemes [7, 10] and the proposals is summarized in Table I.
This paper is organized as follows: In Section II, we review necessary contents from Ogawa et al. [7], a new encoding procedure is given for [7], and we prove that the correspondence between secrets and shares is retained from [7]. Section III has a similar structure to Section II. In Section III, we review [10], propose a new encoding procedure for [10], then we prove that the same correspondence between secrets and shares. Throughout Sections II and III, illustrating examples of the original and the proposed encoding are included. In Section IV, we give concluding remarks and future research agendas.
II Advance sharing with quantum ramp secret sharing scheme in [7]
II-A Short review of quantum secret sharing
In this subsection, we briefly describe the quantum secret sharing considered in this paper, and its security model of secret sharing schemes. In short, it is exactly the same as the paper by Cleve, Gottesman, and Lo [3] and Ogawa et al. [7]. When both secret and shares are classical information, this model is equivalent to Shamir-Blakley scheme [15].
[…]
Let $q$ be a prime power, $\mathbf{F}_{q}$ be the finite field with $q$ elements, and $\{|i\rangle:i\in\mathbf{F}_{q}\}$ be an orthonormal basis of $\mathcal{H}_{q}$. For $\vec{s}=(s_{1}$, …, $s_{L})\in\mathbf{F}_{q}^{L}$ we define
$|\vec{s}\rangle=|s_{1}\rangle\otimes\cdots\otimes|s_{L}\rangle\in\mathcal{H}_{q}^{\otimes L}.$
The dealer encodes the quantum secret $|\vec{s}\rangle$ (or a linear combination of $|\vec{s}\rangle$) into some pure state $|\varphi\rangle\in\mathcal{H}_{q}^{\otimes n}$. The dealer distributes each qudit in $|\varphi\rangle$ to each participant. Observe that no measurement, classical communications, nor teleportation of quantum states are involved here, in contrast to [16]. Also observe that the purpose and the procedure are different from those in quantum secure direct communication (QSDC) [17]. There seems no explicit relationship between the QSDC and the quantum secret sharing considered here.
A set $A\subset\{1$, …, $n\}$ is said to be qualified if $|\vec{s}\rangle$ can be reconstructed from $\mathrm{Tr}_{\overline{A}}[|\varphi\rangle\langle\varphi|]$, said to be forbidden if $\mathrm{Tr}_{\overline{A}}[|\varphi\rangle\langle\varphi|]$ is independent of $|\vec{s}\rangle$, and said to be intermediate if $A$ is neither qualified nor forbidden [7], where $\mathrm{Tr}_{\overline{A}}$ denotes the partial trace over $\overline{A}=\{1$, …, $n\}\setminus A$. As mentioned in Section I, a quantum secret sharing scheme is said to be a ramp scheme if there is no intermediate set.
In Sections II-E and III-D, we will show that our proposed encoding procedures retain the correspondences between quantum secrets and quantum shares from [7, 10]. This means that the sets of qualified sets and forbidden sets remains the same as [7, 10].
More importantly, since the correspondences between secrets and shares are the same as [7, 10], new encoding procedures are as much secure as the original ones under any security models and assumptions, and any new security argument is unnecessary unless a security property not considered in [7, 10] is required.
Although we can consider cheating participants among legitimate ones and adversaries outside of legitimate participants [18, 19], we do not consider these scenarios. We do not consider approximate quantum secret sharing [20], neither, as they are not considered in the original encoding procedures [7, 10]. However, since those other scenarios are also important, advance sharing in them should also be investigated in future.
The standard metrics for evaluating secret sharing schemes are coding rates and access structures [7], and sometimes the strong security [9, 10]. The proposed encoding procedures retains those evaluation metrics from the original schemes [7, 10], and they enable advance sharing that enhances applicability of secret sharing schemes to wider scenarios.
II-B Ogawa et al.’s scheme
Let $n$ be the number of shares/participants, and we assume $n\leq q-1$. Let $\alpha_{1}$, …, $\alpha_{n}$ be distinct nonzero [^1] Ogawa et al. [7] allowed zero but it was a mistake. elements in $\mathbf{F}_{q}$. We will review the construction of a $(k,L,n)$ ramp quantum scheme. For $\vec{c}=(c_{1}$, …, $c_{k})\in\mathbf{F}_{q}^{k}$, we define the polynomial $f_{\vec{c}}(x)$ by
$f_{\vec{c}}(x)=c_{1}+c_{2}x+\cdots+c_{k}x^{k-1}.$ (1)
A quantum secret $|\vec{s}\rangle\in\mathcal{H}_{q}^{\otimes L}$ is encoded to
$\frac{1}{\sqrt{q^{k-L}}}\sum_{\vec{c}\in D(\vec{s})}|f_{\vec{c}}(\alpha_{1})\rangle\otimes|f_{\vec{c}}(\alpha_{2})\rangle\otimes\cdots\otimes|f_{\vec{c}}(\alpha_{n})\rangle,$ (2)
where $D(\vec{s})$ is the set of vectors $\vec{c}\in\mathbf{F}_{q}^{k}$ whose leftmost $L$ components are the same as those in $\vec{s}$. The quantum state in Eq.  (2) consists of $n$ qudits. The $i$-th qudit in Eq.  (2) is distributed to the $i$-th participant.
II-C Example of the original encoding
Let $q=4$, $n=3$, $k=2$, $L=1$, and $\alpha_{i}=\alpha^{i}\in\mathbf{F}_{4}$, where $\alpha$ is a primitive element of $\mathbf{F}_{4}$. The standard encoding procedure in Eq. (2) encodes a quantum secret $|\vec{s}\rangle$ for $\vec{s}=(s)\in\mathbf{F}_{4}^{1}$ into $\frac{1}{\sqrt{4}}\sum_{\vec{c}\in D(\vec{s})}|f_{\vec{c}}(\alpha_{1})\rangle\otimes|f_{\vec{c}}(\alpha_{2})\rangle\otimes|f_{\vec{c}}(\alpha_{3})\rangle=\frac{1}{\sqrt{4}}(|s$, $s$, $s\rangle+|s+\alpha$, $s+\alpha^{2}$, $s+1\rangle+|s+\alpha^{2}$, $s+1$, $s+\alpha\rangle+|s+1$, $s+\alpha$, $s+\alpha^{2}\rangle).$ As every qudit seems to depend on the quantum secret $|\vec{s}\rangle$, from the encoding procedure in Ogawa et al. [7], it seems unclear how one share is advance shareable.
II-D Proposed advance sharing procedure for [7]
We will propose an advance sharing procedure which retains the correspondence between a quantum secret and quantum shares given in Eq. (2), where the $(k,L,n)$ quantum ramp scheme was considered.
We will introduce an elementary lemma in linear algebra and polynomials.
Lemma 1
Let $\mathcal{P}_{m}$ be the set of univariate polynomials $f(x)$ over $\mathbf{F}_{q}$ with $\deg(f)<m$. Consider the evaluation map $\mathrm{ev}(f)=(f(\alpha_{1})$, …, $f(\alpha_{n}))$, where $\alpha_{1}$, …, $\alpha_{n}$ are pairwise distinct as before while $\alpha_{1}$, …, $\alpha_{n}$ may contain $0\in\mathbf{F}_{q}$. The map $\mathrm{ev}$ is a linear map from $\mathcal{P}_{m}$ to $\mathbf{F}_{q}^{n}$. If $m\leq n$ then $\mathrm{ev}$ is injective, and if $m\geq n$ then $\mathrm{ev}$ is surjective.
Proof:
Suppose that $m\leq n$. If $\mathrm{ev}(f)=\vec{0}$ then $f(\alpha_{1})=\cdots=f(\alpha_{n})=0$. Since $\deg(f)<m\leq n$, this means that $f(x)$ is the zero polynomial. We have seen that $\ker(\mathrm{ev})=\{0\}$ when $m\leq n$, which means that $\mathrm{ev}$ is an injective linear map.
By the previous discussion, $\mathrm{ev}$ is injective on $\mathcal{P}_{n}$. Since $\dim\mathcal{P}_{n}=n$, the image $\mathrm{ev}(\mathcal{P}_{n})$ of $\mathcal{P}_{n}$ under the map $\mathrm{ev}$ is $\mathbf{F}_{q}^{n}$. This means that $\mathrm{ev}$ is surjective on $\mathcal{P}_{m}$ if $m\geq n$. ∎
We consider to distribute $k-L$ shares before $|\vec{s}\rangle$ is given to the dealer. By reassigning indices, we may assume that the dealer wants to advance share the 1st to the $(k-L)$-th shares. Let
$|\Psi\rangle=\frac{1}{\sqrt{q^{k-L}}}\sum_{\vec{r}\in\mathbf{F}_{q}^{k-L}}|\vec{r}\rangle\otimes|\vec{r}\rangle,$ (3)
where $|\vec{r}\rangle=|r_{1}\rangle\otimes\cdots\otimes|r_{k-L}\rangle$ for $\vec{r}=(r_{1}$, …, $r_{k-L})\in\mathbf{F}_{q}^{k-L}$. Note that $|\Psi\rangle$ consists of $2(k-L)$ qudits, as $|\vec{r}\rangle$ consists of $(k-L)$ qudits.
Suppose that the quantum secret is $|\vec{s}\rangle=|s_{1}\rangle\otimes\cdots\otimes|s_{L}\rangle\in\mathcal{H}_{q}^{\otimes L}$ with $\vec{s}=(s_{1}$, …, $s_{L})\in\mathbf{F}_{q}^{L}$. For $\vec{r}\in\mathbf{F}_{q}^{k-L}$ and $\vec{s}$, let $g_{\vec{r},\vec{s}}(x)$ be a univariate polynomial such that $g_{\vec{r},\vec{s}}(x)\in\{f_{\vec{c}}(x):\vec{c}\in D(\vec{s})\}$ and $g_{\vec{r},\vec{s}}(x)(\alpha_{i})=r_{i}$ for $i=1$, …, $k-L$.
Proposition 2
For given $\vec{r}$ and $\vec{s}$, the polynomial $g_{\vec{r},\vec{s}}(x)$ exists uniquely.
Proof:
Let $f_{\vec{c}}(x)$ for $\vec{c}\in D(\vec{s})$ as defined in Eqs. (1) and (2). Recall that the coefficients $c_{1}$, …, $c_{L}$ are known as $c_{1}=s_{1}$, …, $c_{L}=s_{L}$ by the condition $\vec{c}\in D(\vec{s})$. We will regard the coefficients $c_{L+1}$, …, $c_{k}$ in $f_{\vec{c}}(x)$ as $k-L$ unknowns in a system of linear equations. For $i=1$, …, $k-L$, the condition $g_{\vec{r},\vec{s}}(x)(\alpha_{i})=r_{i}$ can be written as
$c_{L+1}+c_{L+1}\alpha_{i}+\cdots+c_{k}\alpha_{i}^{k-L-1}$ (4)
$=$ $\frac{r_{i}-s_{1}-s_{2}\alpha_{i}-\cdots-s_{L}\alpha_{i}^{L-1}}{\alpha_{i}^{L}}.$ (5)
The polynomial $f(x)=c_{L+1}+\cdots+c_{k}x^{k-L-1}\in\mathcal{P}_{k-L}$ corresponds to a solution of the system of linear equations if $\mathrm{ev}(f)=((r_{1}-s_{1}-s_{2}\alpha_{1}-\cdots-s_{L}\alpha_{1}^{L-1})/\alpha_{1}^{L}$, $(r_{2}-s_{1}-s_{2}\alpha_{2}-\cdots-s_{L}\alpha_{2}^{L-1})/\alpha_{2}^{L}$, …, $(r_{k-L}-s_{1}-s_{2}\alpha_{k-L}-\cdots-s_{L}\alpha_{k-L}^{L-1})/\alpha_{k-L}^{L})$, where the map $\mathrm{ev}$ is from $\mathcal{P}_{k-L}$ to $\mathbf{F}_{q}^{k-L}$. By Lemma 1 we can see that the solution of the system of linear equations considered here exists uniquely, which ensures unique existence of $g_{\vec{r},\vec{s}}(x)$ for given $\vec{r}$ and $\vec{s}$. ∎
Remark 3
The coefficients of $g_{\vec{r},\vec{s}}(x)$ can be more explicitly described in terms of $\vec{r}$ and $\vec{s}$, as follows. As in the proof of Proposition 2, let $f(x)=c_{L+1}+\cdots+c_{k}x^{k-L-1}$. We also use $(b_{1}$, …, $b_{k-L})=((r_{1}-s_{1}-s_{2}\alpha_{1}-\cdots-s_{L}\alpha_{1}^{L-1})/\alpha_{1}^{L}$, $(r_{2}-s_{1}-s_{2}\alpha_{2}-\cdots-s_{L}\alpha_{2}^{L-1})/\alpha_{2}^{L}$, …, $(r_{k-L}-s_{1}-s_{2}\alpha_{k-L}-\cdots-s_{L}\alpha_{k-L}^{L-1})/\alpha_{k-L}^{L})$. As shown in the proof of Proposition 2, we have $f(\alpha_{i})=b_{i}$ for $i=1$, …, $k-L$. To determine $f(x)$, we can use the Lagrange interpolation formula [15, Chapter 13]. For $i=1,\ldots,k-L$, let
$\ell_{i}(x)=\prod_{1\leq j\leq k-L,j\neq i}\frac{x-\alpha_{j}}{\alpha_{i}-\alpha_{j}}.$
Then $f(x)=c_{L+1}+\cdots+c_{k}x^{k-L-1}$ is given as
$f(x)$ $=$ $\sum_{i=1}^{k-L}b_{i}\ell_{i}(x)$
$=$ $\sum_{i=1}^{k-L}\frac{r_{i}-s_{1}-s_{2}\alpha_{i}-\cdots-s_{L}\alpha_{i}^{L-1}}{\alpha_{i}^{L}}\ell_{i}(x)$
and we have
$g_{\vec{r},\vec{s}}(x)$
$=$ $s_{1}+s_{2}x+\cdots+s_{L}x^{L-1}+c_{L+1}x^{L}+\cdots+c_{k}x^{k-1}$
$=$ $x^{L}f(x)+\sum_{i=1}^{L}s_{i}x^{i-1}.$
Now we are ready to describe a new encoding procedure enabling advance sharing. By Proposition 2, we can define a unitary map $U_{\mathrm{enc}}$ from $\mathcal{H}_{q}^{\otimes k}$ to $\mathcal{H}_{q}^{\otimes(n-k+L)}$ sending a quantum state $|\vec{r}\rangle|\vec{s}\rangle$ to $|g_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|g_{\vec{r},\vec{s}}(\alpha_{n})\rangle$. Note that $k=n-k+L$ by [7, Lemma 2].
The proposed procedure for advance sharing is as follows:
  1. 1.
    The dealer distributes $k-L$ qudits in the left half of Eq. (3).
  2. 2.
    After the quantum secret $|\vec{s}\rangle$ is given, the dealer applies $U_{\mathrm{enc}}$ to the remaining half of Eq. (3) and $|\vec{s}\rangle$. Then the quantum state of all the $n$ shares becomes
    $\frac{1}{\sqrt{q^{k-L}}}\sum_{\vec{r}\in\mathbf{F}_{q}^{k-L}}|\vec{r}\rangle\otimes|g_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|g_{\vec{r},\vec{s}}(\alpha_{n})\rangle.$ (6)
Remark 4
When $k-L-1$ or a fewer shares are advance shared, the dealer can simply keep some shares in the advance sharing phase in our proposal. On the other hand, it is impossible to advance share $k-L+1$ or more shares. When a set $J$ of shares is not forbidden, the shares in $J$ depend on quantum secrets [7, Theorem 2]. Thus, in order for a set $J$ of shares to be advance shareable, $J$ must be a forbidden set. In Ogawa et al.’s $(k,L,n)$ scheme, any $k-L+1$ or more shares cannot form a forbidden set, and cannot be advance shareable. Our proposal makes advance shareable sets as large as possible.
II-E Proof of correctness
It is clear from the last subsection that the proposed procedure can distribute $k-L$ or a fewer shares before a secret is given to the dealer. On the other hand, at this point it is unknown whether or not the proposed encoding procedure produces the same quantum states of shares as the original procedure [7]. In order to show their sameness, in this subsection, we will prove that the proposed procedure gives the same quantum state of $n$ shares as the original scheme by Ogawa et al. [7] for a given quantum secret $|\vec{s}\rangle$. In Eq. (2), the set of indices appearing in the quantum state is
$V_{1}=\{(f_{\vec{c}}(\alpha_{1}),\ldots,f_{\vec{c}}(\alpha_{n})):\vec{c}\in D(\vec{s})\}.$ (7)
In Eq. (6), the set of indices appearing in the quantum state is
$V_{2}$ $=$ $\{(r_{1},\ldots,r_{k-L},g_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,g_{\vec{r},\vec{s}}(\alpha_{n}))$ (8)
$:\vec{r}\in\mathbf{F}_{q}^{k-L}\}.$
Theorem 5
For a fixed quantum secret $|\vec{s}\rangle$ for $\vec{s}\in\mathbf{F}_{q}^{L}$, the proposed procedure gives the same quantum state of shares as the original scheme by Ogawa et al. [7].
Proof:
Equation (2) can be written as
$\frac{1}{\sqrt{q^{k-L}}}\sum_{(v_{1},\ldots,v_{n})\in V_{1}}|v_{1}\rangle\otimes\cdots\otimes|v_{n}\rangle,$ (9)
and Eq. (6) can be written as
$\frac{1}{\sqrt{q^{k-L}}}\sum_{(v_{1},\ldots,v_{n})\in V_{2}}|v_{1}\rangle\otimes\cdots\otimes|v_{n}\rangle.$ (10)
To prove the theorem, we have to prove that Eqs. (9) and (10) are equal. In order to prove the equality between Eqs. (9) and (10), it is enough to show that the sets (7) and (8) are the same. To show the equality between two sets (7) and (8), in Eq. (11) we will show that there is a one-to-one correspondence between elements in the two sets (7) and (8).
For fixed $\vec{s}\in\mathbf{F}_{q}^{L}$ and $\vec{r}\in\mathbf{F}_{q}^{k-L}$, by Proposition 2 there exists a unique polynomial $f_{\vec{c}}$ with $\vec{c}\in D(\vec{s})$. This correspondence gives a bijection sending $\vec{r}\in\mathbf{F}_{q}^{k-L}$ to $f_{\vec{c}}$ with $\vec{c}\in D(\vec{s})$. By the definition of $g_{\vec{r},\vec{s}}$, we have
$(r_{1},\ldots,r_{k-L},g_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,g_{\vec{r},\vec{s}}(\alpha_{n}))\in V_{2}$ (11)
$=$ $(g_{\vec{r},\vec{s}}(\alpha_{1}),\ldots,g_{\vec{r},\vec{s}}(\alpha_{n}))$
$=$ $(f_{\vec{c}}(\alpha_{1}),\ldots,f_{\vec{c}}(\alpha_{n}))\in V_{1},$
which shows the theorem, where $\vec{c}$ is chosen according to $\vec{r}$. ∎
II-F Example of the proposed encoding
We reuse definitions from Section II-C. Let $q=4$, $n=3$, $k=2$, $L=1$, and $\alpha_{i}=\alpha^{i}\in\mathbf{F}_{4}$, where $\alpha$ is a primitive element of $\mathbf{F}_{4}$. Note that since $q$ is not a prime number, this case cannot be handled in the previous research [14]. In our proposal, the dealer prepares 2-qudit entangled state $\frac{1}{2}\sum_{r\in\mathbf{F}_{4}}|r\rangle|r\rangle$, and send 1 qudit in it. Then the dealer is given a quantum secret $|\vec{s}\rangle$ for $\vec{s}=(s)\in\mathbf{F}_{4}^{1}$. Then we have
$g_{r,s}(x)=\frac{r-s}{\alpha_{1}}x+s.$
The unitary map $U_{\mathrm{enc}}$ sends $|r\rangle|s\rangle$ to $|\frac{r-s}{\alpha_{1}}\alpha_{2}+s\rangle|\frac{r-s}{\alpha_{1}}\alpha_{3}+s\rangle$. Now it is clear that the left most share can be distributed before the secret $|s\rangle$ is given, in contrast to Section II-C.
III Advance sharing with quantum ramp secret sharing scheme in [10]
III-A Zhang and Matsumoto’s scheme
As mentioned before, Zhang and Matsumoto [10] also proposed a $(k,L,n)$ ramp quantum secret sharing scheme. Their proposal has strong security in contrast to Ogawa et al. [7], and the maximum possible number of participants was smaller, that is, $n\leq q-L$. As in Section II, $\alpha_{1}$, …, $\alpha_{n}$ are pairwise distinct elements in $\mathbf{F}_{q}$. We choose extra elements $\beta_{1}$, …, $\beta_{L}$ such that all $\alpha_{1}$, …, $\alpha_{n}$ and $\beta_{1}$, …, $\beta_{L}$ are pairwise distinct. In [10] any of $\alpha_{1}$, …, $\alpha_{n}$ and $\beta_{1}$, …, $\beta_{L}$ can be $0\in\mathbf{F}_{q}$. For $\vec{s}\in\mathbf{F}_{q}^{L}$, we define
$D_{\mathrm{ZM}}(\vec{s})=\{\vec{c}\in\mathbf{F}_{q}^{k}:f_{\vec{c}}(\beta_{i})=s_{i}\mbox{ for }i=1,\ldots,L\},$ (12)
where $f_{\vec{c}}(x)$ is as defined in Eq. (1). For a given quantum secret $|\vec{s}\rangle$, in [10] the quantum state of $n$ shares are obtained as
$\frac{1}{\sqrt{q^{k-L}}}\sum_{\vec{c}\in D_{\mathrm{ZM}}(\vec{s})}|f_{\vec{c}}(\alpha_{1})\rangle\otimes|f_{\vec{c}}(\alpha_{2})\rangle\otimes\cdots\otimes|f_{\vec{c}}(\alpha_{n})\rangle,$ (13)
III-B Example of the original encoding
Let $q=7$, $n=4$, $k=3$, $L=2$, $(\alpha_{1}$, …, $\alpha_{4})=(6$, $2$, $4$, 5)5)5 ) and $(\beta_{1}$, $\beta_{2})=(1$, 3)3)3 ). This example is the same as [10, Example 2]. For $\vec{s}=(s_{1}$, $s_{2})$, $D_{\mathrm{ZM}}(\vec{s})$ contains $(c_{1}$, …, $c_{3})$ if and only if
$\left\{\begin{array}[]{rcl}c_{1}+c_{2}+c_{3}&=&s_{1},\\ c_{1}+3c_{2}+2c_{3}&=&s_{2}.\end{array}\right.$
The solution of the above system of linear equations for $(c_{1}$, …, $c_{3})$ is $D_{\mathrm{ZM}}(\vec{s})=\{(5s_{1}+3s_{2}+3c_{3},3s_{1}+4s_{2}+3c_{3},c_{3}):c_{3}\in\mathbf{F}_{7}\}$, which is the same [10, Eq. (3)]. For a given quantum state $|\vec{s}\rangle=|s_{1}\rangle\otimes|s_{2}\rangle$, the quantum state of $4$ shares is (the normalizing constant $1/\sqrt{7}$ is omitted) $\sum_{c_{3}=0}^{6}\otimes_{i=1}^{4}|\alpha_{i}^{2}c_{3}+\alpha_{i}(3s_{1}+4s_{2}+3c_{3})+(5s_{1}+3s_{2}+3c_{3})\rangle=|2s_{1}+6s_{2}\rangle\otimes|4s_{1}+4s_{2}\rangle\otimes|3s_{1}+5s_{2}\rangle\otimes|6s_{1}+2s_{2}\rangle+\cdots$, which is equivalent to [10, Eq. (5)]. Since the quantum state of every qudit seems to depend on $s_{1}$ and $s_{2}$, from the encoding procedure in [10], it seems unclear how one share is advance shareable.
III-C Proposed advance sharing procedure for [10]
We consider to distribute $k-L$ shares before $|\vec{s}\rangle$ is given to the dealer. By reassigning indices, we may assume that the dealer wants to advance share the 1st to the $(k-L)$-th shares. Let $|\Psi\rangle$ as defined in Eq. (3).
Suppose that the quantum secret is $|\vec{s}\rangle=|s_{1}\rangle\otimes\cdots\otimes|s_{L}\rangle\in\mathcal{H}_{q}^{\otimes L}$ with $\vec{s}=(s_{1}$, …, $s_{L})\in\mathbf{F}_{q}^{L}$. For $\vec{r}\in\mathbf{F}_{q}^{k-L}$ and $\vec{s}$, let $h_{\vec{r},\vec{s}}(x)$ be a univariate polynomial such that $h_{\vec{r},\vec{s}}(x)\in\{f_{\vec{c}}(x)$ : $\vec{c}\in D_{\mathrm{ZM}}(\vec{s})\}$ and $h_{\vec{r},\vec{s}}(x)(\alpha_{i})=r_{i}$ for $i=1$, …, $k-L$.
Proposition 6
For given $\vec{r}$ and $\vec{s}$, the polynomial $h_{\vec{r},\vec{s}}(x)$ exists uniquely.
Proof:
Let $\mathcal{P}_{k}$ as defined in Lemma 1, and consider a polynomial $f(x)\in\mathcal{P}_{k}$. The condition $f(x)\in\{f_{\vec{c}}(x)$ : $\vec{c}\in D_{\mathrm{ZM}}(\vec{s})\}$ means that $f(\beta_{i})=s_{i}$ for $i=1$, …, $L$. Together with the conditions $f(\alpha_{i})=r_{i}$ for $i=1$, …, $k-L$, the required $h_{\vec{r},\vec{s}}(x)$ is a polynomial $f(x)\in\mathcal{P}_{k}$ such that $\mathrm{ev}^{\prime}(f)=(s_{1}$, …, $s_{L}$, $r_{1}$, …, $r_{k-L})$, where $\mathrm{ev}^{\prime}(f)=(f(\beta_{1})$, …, $f(\beta_{L})$, $f(\alpha_{1})$, …, $f(\alpha_{k-L}))$. By Lemma 1, we can see that such a polynomial $f(x)$ exists uniquely, which is $h_{\vec{r},\vec{s}}(x)$. ∎
Remark 7
Similarly to Remark 3, we will give a more explict description of $h_{\vec{r},\vec{s}}(x)$, which is determined by the equality condition $(h_{\vec{r},\vec{s}}(\beta_{1})$, …, $h_{\vec{r},\vec{s}}(\beta_{L})$, $h_{\vec{r},\vec{s}}(\alpha_{1})$, …, $h_{\vec{r},\vec{s}}(\alpha_{k-L}))=(s_{1}$, …, $s_{L}$, $r_{1}$, …, $r_{k-L})$. The determination of $h_{\vec{r},\vec{s}}(x)$ is known as the Lagrange interpolation [15, Chapter 13], which can be computed as follows. In order to use the Lagrange interpolation formula, define $\beta_{L+j}=\alpha_{j}$ for $j=1$, …, $k-L$, and
$\ell_{i}(x)=\prod_{1\leq j\leq k,j\neq i}\frac{x-\beta_{j}}{\beta_{i}-\beta_{j}}.$
Now $h_{\vec{r},\vec{s}}(x)$ can be obtained as
$h_{\vec{r},\vec{s}}(x)=\sum_{i=1}^{L}s_{i}\ell_{i}(x)+\sum_{i=1}^{k-L}r_{i}\ell_{i+L}(x).$
Let us describe a new encoding procedure enabling advance sharing. By Proposition 6, we can define a unitary map $U_{\mathrm{ZM,enc}}$ from $\mathcal{H}_{q}^{\otimes k}$ to $\mathcal{H}_{q}^{\otimes(n-k+L)}$ sending a quantum state $|\vec{r}\rangle|\vec{s}\rangle$ to $|h_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|h_{\vec{r},\vec{s}}(\alpha_{n})\rangle$. Note that $k=n-k+L$ by [7, Lemma 2].
The proposed procedure for advance sharing is as follows:
  1. 1.
    The dealer distribute $k-L$ qudits in the left half of Eq. (3).
  2. 2.
    After the quantum secret $|\vec{s}\rangle$ is given, the dealer applies $U_{\mathrm{ZM,enc}}$ to the remaining half of Eq. (3) and $|\vec{s}\rangle$. Then the quantum state of all the $n$ shares becomes
    $\frac{1}{\sqrt{q^{k-L}}}\sum_{\vec{r}\in\mathbf{F}_{q}^{k-L}}|\vec{r}\rangle\otimes|h_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|h_{\vec{r},\vec{s}}(\alpha_{n})\rangle.$ (14)
Remark 8
When $k-L-1$ or a fewer shares are advance shared, the dealer can simply keep some shares in the advance sharing phase in our proposal as in Remark 4. On the other hand, as in Remark 4, it is impossible to advance share $k-L+1$ or more shares, because in Zhang and Matsumoto’s $(k,L,n)$ scheme [10], any $k-L+1$ or more shares cannot form a forbidden set. Therefore, similarly to Remark 4, our proposal makes advance shareable sets as large as possible.
III-D Proof of correctness
By the same reason explained at the beginning of Section II-E, in this subsection, we will prove that the proposed procedure gives the same quantum state of shares as the original [10] for a given quantum secret $|\vec{s}\rangle$. In Eq. (13), the set of indices appearing in the quantum state is
$\{(f_{\vec{c}}(\alpha_{1}),\ldots,f_{\vec{c}}(\alpha_{n})):\vec{c}\in D_{\mathrm{ZM}}(\vec{s})\}.$ (15)
In Eq. (14), the set of indices appearing in the quantum state is
$\{(r_{1},\ldots,r_{k-L},h_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,h_{\vec{r},\vec{s}}(\alpha_{n}):\vec{r}\in\mathbf{F}_{q}^{k-L}\}.$ (16)
Theorem 9
For a fixed quantum secret $|\vec{s}\rangle$ for $\vec{s}\in\mathbf{F}_{q}^{L}$, the proposed procedure gives the same quantum state of shares as the original [10].
Proof:
In order to prove the theorem, it is enough to show that the sets (15) and (16) are the same. For fixed $\vec{s}\in\mathbf{F}_{q}^{L}$ and $\vec{r}\in\mathbf{F}_{q}^{k-L}$, by Proposition 6 there exists a unique polynomial $f_{\vec{c}}(x)$ with $\vec{c}\in D_{\mathrm{ZM}}(\vec{s})$. This correspondence gives a bijection between $\vec{r}\in\mathbf{F}_{q}^{k-L}$ to $f_{\vec{c}}(x)$ with $\vec{c}\in D(\vec{s})$. By the definition of $h_{\vec{r},\vec{s}}$, we have
$(r_{1},\ldots,r_{k-L},h_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,h_{\vec{r},\vec{s}}(\alpha_{n}))$
$=$ $(h_{\vec{r},\vec{s}}(\alpha_{1}),\ldots,h_{\vec{r},\vec{s}}(\alpha_{n}))$
$=$ $(f_{\vec{c}}(\alpha_{1}),\ldots,f_{\vec{c}}(\alpha_{n})),$
which shows the theorem, where $\vec{c}$ is chosen according to $\vec{r}$. ∎
III-E Example of the proposed encoding
We reuse definitions from Section III-B. Let $q=7$, $n=4$, $k=3$, $L=2$, $(\alpha_{1}$, …, $\alpha_{4})=(6$, $2$, $4$, 5)5)5 ) and $(\beta_{1}$, $\beta_{2})=(1$, 3)3)3 ). In our proposal, the dealer prepares 2-qudit entangled state $\frac{1}{\sqrt{7}}\sum_{r\in\mathbf{F}_{7}}|r\rangle|r\rangle$, and send 1 qudit in it. Then the dealer is given a quantum secret $|\vec{s}\rangle$ for $\vec{s}=(s_{1},s_{2})\in\mathbf{F}_{7}^{2}$. Then we have
$h_{r,s_{1},s_{2}}(x)=(r-2s_{1}+s_{2})x^{2}+(4s_{1}-4r)x+(3r-s_{1}-s_{2}),$
because $h_{r,s_{1},s_{2}}(x)$ satisfies $h_{r,s_{1},s_{2}}(\alpha_{1})=h_{r,s_{1},s_{2}}(6)=r$ , $h_{r,s_{1},s_{2}}(\beta_{1})=h_{r,s_{1},s_{2}}(1)=s_{1}$, and $h_{r,s_{1},s_{2}}(\beta_{2})=h_{r,s_{1},s_{2}}(3)=s_{2}$. The unitary map $U_{\mathrm{ZM,enc}}$ sends $|r\rangle|s_{1}\rangle|s_{2}\rangle$ to $|h_{r,s_{1},s_{2}}(\alpha_{2})\rangle|h_{r,s_{1},s_{2}}(\alpha_{3})\rangle|h_{r,s_{1},s_{2}}(\alpha_{3})\rangle$. Now it is clear that the left most share can be distributed before the secret $|\vec{s}\rangle$ is given, in contrast to Section III-B.
IV Concluding remarks
In this paper, we proposed new encoding procedures for Ogawa et al.’s [7] and Zhang-Matsumoto [10] ramp quantum secret sharing schemes, and allowed the dealer to distribute shares in those schemes before secrets are given. This enhances applicability of the secret sharing schemes by Ogawa et al. [7] and by Zhang and Matsumoto [10] to wider scenarios in which some participants are unavailable when the dealer obtains a secret. We also proved that the new proposed encoding procedures retain desirable properties, such as coding rates, access structures, strong security from the original schemes. We also proved that the advance shareable sets are made as large as possible in the proposed encoding procedures.
Similarly to [13], there could exist a general method to construct advance sharing procedure for any stabilizer-based quantum secret sharing, while retaining the correspondence between secrets and shares. Its investigation could be a future research agenda.
As mentioned in Section II-A, it seems also important to study advance sharing with other scenarios of quantum secret sharing, for example, [18, 19, 20]. It is also another research agenda.
Acknowledgment
The author would like to deeply thank reviewers’ helpful comments that improved this paper.